Below is the abbreviated cash budget for XYZ Pty Ltd.

Beginning Cash Balance 25,000
Estimated Cash Receipts
Cash Available 365,000
Estimated Cash Payments
Purchases 150,000
Operating Expenses
Office Equipment Purchases 25,000
Total Estimated Cash Payments
Ending Cash Balance 150,000


(a) Calculate the amount of Estimated Cash Receipts, Operating Expenses, and Total   Estimated Cash Payments.  
(b) What action if any should XYZ Pty Ltd take as a result of preparing this budget? Briefly justify your answer.
Answer #1

SOLUTION

(A)  Estimated Cash Receipts = Total cash available - Beginning cash balance

= 365,000 - 25,000 = 340,000

Total Estimated Cash Payments = Total cash available - Ending cash balance

= 365,000 - 150,000 = 215,000

Operating expenses = Total Estimated Cash Payments - Purchases - Office Equipment Purchases

= 215,000 - 150,000 - 25,000 = 40,000

(B) XYZ Pty Ltd. has made the correct budget as they are not required to take any loan to maintain the ending cash balance.